### v4

# Linear Model of Aggregated Homogeneous Energy Storage Elements with Realizable Dispatch Guarantees

Abstract
To optimize the dispatch of batteries, a model is required that can predict the state of charge (SOC) trajectory for a chosen open-loop power schedule to ensure admissibility (i.e., that schedule can be realized). However, battery dispatch optimization is inherently challenging since batteries cannot simultaneously charge and discharge, which begets a non-convex complementarity constraint. In this paper, we develop a novel composition of energy storage elements that can charge or discharge independently and provide a sufficient linear energy storage model of the composite battery. This permits convex optimization of the composite battery SOC trajectory while guaranteeing admissibility of the resulting (aggregated) power schedule and its disaggregation to the individual energy storage elements.
Index Terms:
Battery, energy storage, optimization, convex, priority-based control policy, complementarity conditions.
I Introduction
Battery energy storage systems (BESS) are increasingly deployed to help facilitate the grid’s uptake of variable renewable energy and improve reliability  [Battery_review_paper, BESS_simultaneous_participation, BESS_renewables]. In many applications, BESS dispatch involves coordinating $N$ battery racks/modules that together comprise a single BESS unit (e.g., DC-coupled elements), $N$ BESS container units in a utility-scale battery storage facility (e.g., AC-coupled elements), or $N$ aggregated BESS units coordinated as a virtual battery/power plant (VPP). However, when optimizing BESS dispatch, the individual elements that make up the BESS are often lumped together into a single composite energy storage asset. The composite dispatch is then optimized and the resulting power schedule is shared equally among the $N$ battery elements. An optimal composite dispatch represents a non-convex problem since the complementarity constraint is formulated with a non-linear constraint or by the use of binary variables [MILP_for_CC, prat_2024].
To obtain a convex relaxation of the dispatch problem, the complementarity constraint is typically omitted [prat_2024]. The resulting power trajectories can then exhibit simultaneous charging and discharging, which cannot be implemented. To recover near-term feasibility, only the net power trajectory is applied. As a result, the state of charge (SOC) becomes a provable lower bound on the SOC when the schedule is realized [RobustNawaf, almassalkhi_model-predictive_2015]. This can lead to unexpected SOC saturation, resulting in an un-realizable power trajectory. Extensive research has investigated the conditions under which the convex relaxation’s state of charge (SOC) coincides with the actual realized SOC. Authors in [K.Baker_DCOPF_No_CC] avoid the use of the complementarity constraint when applied to DCOPF but add a penalty term to the objective function to ensure the exactness of the relaxed model. Ex-post sufficient conditions under which the relaxation becomes exact for an economic dispatch problem are provided in [Sufficient_Cond_for_Relax, Extended_Sufficient_Cond_for_Relax]. Ex-ante conditions for exactness of the relaxed problem are presented in [Lin_Exact_Cond_Ex_Ante], assuming that certain constraints on battery power limits, as well as initial and maximum energy capacities, are met. Furthermore, [positive_prices, Jakob_Pos_Quadratic_Cost, prat_2024] shows that for economic dispatch problems with positive linear and quadratic cost functions, the convex relaxation at optimality would be exact with no simultaneous charging and discharging.
However, the exactness of the relaxed model does not hold in general which may lead to battery saturation and inadmissible power trajectories. A simple example of this is the case with negative prices as shown in [Baldick_Neg_Price_Bad]. For a single time step, this problem can be alleviated by developing tighter relaxations [D.Pozo_paper, elgersma2024], which can also heuristically be applied to time series, albeit without realizability guarantees.
Figure 1: Comparing different composite models after implementing an open-loop predicted dispatch schedule. The relaxed model dispatch is inadmissible, while the robust linear model is admissible but suboptimal. However, this paper’s model achieves near-optimal performance and is admissible.
Additionally, authors in  [RobustNawaf, Mazen_CDC_paper], an alternative robust linear battery model is presented which is a convex inner approximation of the non-convex formulation that sacrifices optimality for guaranteed admissibility. The robust linear model omits the complementarity constraint but envelopes the actual SOC trajectories to guarantee that any feasible power trajectory is realizable. However, the envelopes beget conservativeness as a function of efficiencies and prediction horizon, which limits practicality over longer horizons. In summary, previous work in energy storage optimization that circumvents the complementarity condition either has been application-specific and holds only if specific criteria on the objective function and/or the battery parameters are met, or presents convex restrictions that may yield significantly sub-optimal solutions. Furthermore, previous work has generally (and often, implicitly) assumed an equal power-sharing policy, which limits how individual elements can respond under dynamic conditions.
Thus, this manuscript revisits the modeling of BESS aggregations by presenting a novel composite model $N$ identical and independently controllable battery elements, where simultaneous charging and discharging power trajectories at the composite level can be implemented by inhomogeneous disaggregation to the individual elements. The main contributions of the paper are as follows:
  1. 1.
    We present a novel linear model that provides a sufficient representation of aggregations of homogeneous battery storage elements.
  2. 2.
    We develop a priority-based strategy, based on element SOC, to perform the composite-to-individual disaggregation.
  3. 3.
    The proposed model guarantees that all individual element-wise power limits, energy limits, and complementarity constraints will be satisfied after disaggregation.
Figure 1 illustrates the open-loop implementation of the dispatch schedules obtained from the proposed model, the relaxed model, and the robust linear model presented in [RobustNawaf] when used to optimize composite battery dispatch to track a given power reference signal. The relaxed model dispatch causes the battery to saturate leading to poor tracking performance. The robust linear model dispatch is realizable; however, yields sub-optimal tracking performance. The proposed model yields near-optimal results and guarantees realizable power trajectories under an unequal power-sharing policy.
The rest of the paper is organized as follows. Section II presents the composite battery model and the associated constraints. Section III presents the proposed disaggregation strategy and the sufficient conditions for the admissibility of the composite power trajectory. Section IV presents simulation results to demonstrate the performance of the proposed model and disaggregation strategy. Finally, Section V concludes the paper.
II Composite Battery Model
In this section, we will present element-wise and composite models of BESS systems. We consider BESS assets composed of an aggregated generic energy storage asset. An example of such a system is shown in Fig 2. It is composed of $N$ identical elements.
Figure 2: Example of a composite battery with AC-coupled battery elements.
For element $i\in\{1,\ldots,N\}$, the charge/discharge power is set in controller time steps $l$ with duration $\delta t$. The charge/discharge power between $l$ and $l+1$, the charge/discharge efficiency and the SOC at timestep $l$ are denoted by $P^{i}_{\text{c/d}}(l)$, $\eta_{\text{c/d}}$, and $E^{i}(l)$, respectively. A realizable BESS dispatch must satisfy power limits (${P}_{\text{c/d,max}}$), energy limits (${E}_{\text{max}}$), and complementarity constraints for all battery elements ($\forall i=1,\ldots,N$) and time steps ($\forall l=0,\ldots,L-1$). Note that all elements have identical power and energy limits. This results in (1), where $E^{i}_{0}$ is the initial SOC of battery element $i$:
$E^{i}(l+1)=E^{i}(l)+\delta t\eta_{\text{c}}P_{\text{c}}^{i}(l)-\delta t\frac{1}{\eta_{\text{d}}}P_{\text{d}}^{i}(l)$ (1a)
$E^{i}(0)=E^{i}_{0}$ (1b)
$0\leq E^{i}(l+1)\leq{E}_{\text{max}}$ (1c)
$0\leq P_{\text{c/d}}^{i}(l)\leq{P}_{\text{c/d,max}}$ (1d)
$0=P_{\text{c}}^{i}(l)P_{\text{d}}^{i}(l).$ (1e)
The scheduling of the composite battery is done with (optionally) longer scheduler time steps of duration $\Delta t=M\delta t,M\in\mathbb{N}^{+}$. We denote these scheduling time steps with index $[k]$ in square brackets, where $k=0,\ldots,K-1$ and $L=MK$. Constant power in step $k$ implies the following constraint for all $l,k$ with $k=\lfloor\frac{l}{M}\rfloor$:
$\sum_{i=1}^{N}P_{\text{c}}^{i}(l)=P_{\text{c}}[k]\quad\land\quad\sum_{i=1}^{N}P_{\text{d}}^{i}(l)=P_{\text{d}}[k].$ (2)
Typically, it is assumed that a composite BESS system is modeled and operated under balanced conditions: the dispatcher (e.g., EMS) maintains equal SOC ($E_{0}^{i}=E_{0}~{}~{}\forall i$) and adopts an equal power sharing policy across its elements. This results in a composite battery with power and energy limits of $N{P}_{\text{c/d,max}}$ and $N{E}_{\text{max}}$ as shown in (3):
$E[k+1]$ $=E[k]+\Delta t\eta_{\text{c}}P_{\text{c}}[k]-\Delta t\frac{1}{\eta_{\text{d}}}P_{\text{d}}[k]$ (3a)
$E[0]$ $=\sum_{i}E^{i}_{0}$ (3b)
00 0 $\leq E[k+1]\leq N{E}_{\text{max}}$ (3c)
00 0 $\leq P_{\text{c/d}}[k]\leq N{P}_{\text{c/d,max}}$ (3d)
00 0 $=P_{\text{c}}[k]P_{\text{d}}[k].$ (3e)
Under the equal power sharing policy, a single complementarity constraint (3e) at the composite battery level is sufficient to satisfy (1e) for all $i$ and all $l$. Under balanced operation, a feasible composite dispatch based on (3) satisfies all elements’ constraints in (1).
The complementarity constraint is often neglected to obtain a convex relaxation of the problem. For example, as
[…]
$\frac{P_{\text{c}}[k]}{N{P}_{\text{c,max}}}$ $+\frac{P_{\text{d}}[k]}{N{P}_{\text{d,max}}}\leq 1,$ (4b)
where the cutting plane in (4b) has been added to tighten the relaxation of the problem. However, such relaxations can result in an optimal power schedule with simultaneous charging and discharging, which cannot be implemented under equal load sharing. Tighter relaxations have been developed [D.Pozo_paper, elgersma2024] to reduce the prevalence of such solutions in real-world problems.
Nevertheless, there are cases where simultaneous charging and discharging (which effectively ‘discards’ energy if $\eta_{c}\eta_{d}<1$) is a desirable outcome of a battery dispatch schedule. For example, this may be the case when energy prices are negative, when the storage unit tracks a dispatch signal, or as byproduct of minimum up- or downtime constraints in a centrally dispatched system.
Such cases can be addressed by a composite battery if we omit the equal load sharing requirement. This way, we can realize simultaneous charging and discharging dispatch schedules by appropriately disaggregating them across the $N$ elements while satisfying (1e). This increases the capability of the composite battery at the cost of computational complexity with $N$ complementary constraints.
In the remainder of this paper, we address this challenge by presenting a sufficient (but not unique) priority-based disaggregation strategy that engenders a linear set of provably realizable composite power trajectories. These trajectories permit simultaneous charging and discharging of the BESS and approximate the full solution space arbitrarily closely.
III Composite Trajectory Disaggregation
A control policy is used to disaggregate a desired piecewise constant composite power inputs $P_{\text{c/d}}[k]$ into $N$ individual element charge/discharge inputs, $\{P^{i}_{\text{c/d}}(l)\}_{i=1}^{N}$. One such controller is the presented priority stack controller discussed below. Sufficient conditions are developed that – in combination with the controller – enable the characterization of $P_{\text{c}}[k]$ and $P_{\text{d}}[k]$ that guarantees (1) is satisfied.
III-A Priority Stack Controller (PSC)
For every controller time-step $l$, the priority stack controller (PSC) sorts all $N$ elements based on the SOC. The element with the lowest SOC has the highest charging priority while the element with the highest SOC has the highest discharging priority. When selecting the highest-priority element for charging/discharging that element’s power is maximized, i.e., $P^{i}_{\text{c/d}}(l)=P_{\text{c/d,max}}$, before actuating another charging/discharging element from the stack. This process is repeated every time-step $l$, adapting to changing element SOCs. The priority stack results in $N_{\text{c}}$ elements with lowest SOCs charging and $N_{\text{d}}$ elements with highest SOCs discharging, as follows:
$N_{\text{c}}(l):=\Big{\lceil}\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\Big{\rceil}\quad\land\quad N_{\text{d}}(l):=\Big{\lceil}\frac{P_{\text{d}}[k]}{{P}_{\text{d,max}}}\Big{\rceil}.$ (5)
The PSC is summarized in Algorithm 1 and begets Property 1.
Algorithm 1 Priority Stack Controller (PSC)
1:Given $P_{\text{c}}[k]$ and $P_{\text{d}}[k]$
2:for each timestep $l$ do
3:    Sort elements based on $E^{i}(l)$ in ascending order
4:    Calculate $N_{\text{c/d}}(l):=\lceil\frac{P_{\text{c/d}}[k]}{P_{\text{c/d,max}}}\rceil$
5:    Assign $P^{i}_{\text{c/d}}(l)$ for first $N_{\text{c}}(l)$/last $N_{\text{d}}(l)$ elements
6:    Update SOC for all elements: $E^{i}(l)\rightarrow E^{i}(l+1)$
7:end for
Property 1.
Under PSC, if elements $i,j$ satisfy $E^{i}(l)\geq E^{j}(l)$ , then $P^{i}_{\text{c}}(l)\leq P^{j}_{\text{c}}(l)$ and $P^{i}_{\text{d}}(l)\geq P^{j}_{\text{d}}(l)$.
III-B Guaranteeing Element Complementarity Constraints
Next, we analyze the effects of the PSC on (1), on complementarity conditions (1e), and on energy limits (1c). To ensure the complementarity constraints are satisfied at the element level, $(P_{\text{c}}^{i}(l)P_{\text{d}}^{i}(l)=0\quad\forall i,\forall l)$, the priority stack cannot issue a charging and discharging demand to the same battery element $i$ simultaneously. Thus, any overlap in the selection of $N_{\text{c}}(l)$ charging and $N_{\text{d}}(l)$ discharging elements imply a violation of the complementarity constraints, which must, therefore, satisfy
$N_{\text{c}}(l)+N_{\text{d}}(l)\leq N,\quad\forall l.$ (6)
Proposition 1.
Under the PSC, if $P_{\text{c}}[k]$ and $P_{\text{d}}[k]$ satisfy
$\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}+\frac{P_{\text{d}}[k]}{{P}_{\text{d,max}}}\leq N-1,$ (7)
then the resulting $N_{\text{c}}(l)$ and $N_{\text{d}}(l)$ satisfy (6).
Proof.
Define $\Gamma_{\text{c}}:=\lfloor\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\rfloor$ and $\Delta_{\text{c}}:=\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}-\Gamma_{\text{c}}\in[0,1)$. Then $\Big{\lceil}\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\Big{\rceil}\leq\Gamma_{c}+1$ and, from (7), we get
$\Big{\lceil}\frac{P_{\text{d}}[k]}{{P}_{\text{d,max}}}\Big{\rceil}\leq\Big{\lceil}N-1-(\Gamma_{\text{c}}+\Delta_{\text{c}})\Big{\rceil}=N-1-\Gamma_{\text{c}}.$ (8)
Adding the inequalities, we get $N_{c}(l)+N_{d}(l)\leq N$ $\forall l$. ∎
Thus, adding (7) to the composite formulation in (3) will satisfy element complementarity constraints under PSC. The feasible regions dictated by (6) and (7) are shown in Fig. 3.
Figure 3: Feasible region for $N=10$ and $P_{\text{c/d,max}}=0.5$MW (left) and $N=30$ and $P_{\text{c/d,max}}=0.17$MW (right). SOC limits are omitted here.
III-C Relating Element and Composite Energy Limit Constraints
Similarly, we want to relate element energy limits to those in the composite formulation. Under PSC, the battery elements are charged/discharged individually every $l$, which results in differing SOC trajectories. The spread of element SOCs can impact realizability of the composite dispatch when individual elements hit their SOC limits. Constraint violations can be avoided by quantifying the maximum possible SOC spread under PSC and using that as an energy buffer applied to the composite battery energy limits.
Proposition 2.
Under the PSC, a composite dispatch that satisfies
$N\varepsilon\leq E[k]\leq N(E_{\text{max}}-\varepsilon),$ (9)
with
$\varepsilon:=\delta t\left(\eta_{\text{c}}{P}_{\text{c,max}}+\frac{{P}_{\text{d,max}}}{\eta_{\text{d}}}\right),$ (10)
will not violate any individual element’s energy limit, provided that (i) $\max_{i,j}\{|E^{i}_{0}-E^{j}_{0}|\}\leq\varepsilon$ and (ii) the control time step $\delta t=\Delta t/M$ is sufficiently small to ensure $\varepsilon\leq\frac{1}{2}E_{\text{max}}$.
Proof.
Define the maximum difference in SOC across all elements as $\Delta E(l):=\max_{i,j}\{|E^{i}(l)-E^{j}(l)|\}$ (i.e., the SOC spread). Consider element SOCs $E^{i}(l)$ and $E^{j}(l)$ and, without loss of generality, let $E^{i}(l)\geq E^{j}(l)$. Using (1a) the following is obtained,
$E^{i}(l+1)-E^{j}(l+1)$ $=\left(E^{i}(l)-E^{j}(l)\right)+$ (11)
$\frac{\Delta t}{M}\eta_{\text{c}}\left(P^{i}_{\text{c}}(l)-P^{j}_{\text{c}}(l)\right)$ $-\frac{\Delta t}{M\eta_{\text{d}}}\left(P^{i}_{\text{d}}(l)-P^{j}_{\text{d}}(l)\right).$
From Property 1 and (1d), we have $-P_{\text{c,max}}\leq P^{i}_{\text{c}}(l)-P^{j}_{\text{c}}(l)\leq 0$ and $0\leq P^{i}_{\text{d}}(l)-P^{j}_{\text{d}}(l)\leq P_{\text{d,max}}$, which implies
$E^{i}(l+1)-E^{j}(l+1)\leq$ $E^{i}(l)-E^{j}(l)\leq\Delta E(l),$ (12a)
$E^{i}(l+1)-E^{j}(l+1)\geq$ $-\frac{\Delta t}{M}\left(\eta_{\text{c}}P_{\text{c,max}}+\frac{P_{\text{d,max}}}{\eta_{\text{d}}}\right)=-\varepsilon,$ (12b)
where the inequality in (12b) uses $E^{i}(l)-E^{j}(l)\geq 0$. Considering all $i,j$, we can bound $\Delta E(l+1)\leq\max\left\{\Delta E(l),\varepsilon\right\}$. Moreover, we see that $\Delta E(l)\leq\varepsilon\Rightarrow\Delta E(l+1)\leq\varepsilon$, so it follows that for a composite battery with initial SOC spread $\Delta E(0)\leq\varepsilon$ (condition (i)), we maintain $\Delta E(l)\leq\varepsilon,\forall l$.
Because energy linearly increases/decreases within each scheduling time step $k$ under PSC, (9) implies its equivalent for controller time steps $l$: $N\varepsilon\leq E(l)\leq N(E_{\text{max}}-\varepsilon)$. Dividing by $N$ begets $\varepsilon\leq\overline{E}(l)\leq E_{\text{max}}-\varepsilon$, where $\overline{E}(l):=\sum_{j}E^{j}(l)/N$ is the average element SOC. Because the SOC spread $\Delta E(l)\leq\varepsilon$, we must have $\overline{E}(l)-\varepsilon\leq E^{i}(l)\leq\overline{E}(l)+\varepsilon,\forall i$. Combined, these inequalities show that (1c) holds for all $i$ and all $l$.
Lastly, we require condition (ii) ($\varepsilon\leq\frac{1}{2}E_{\text{max}}$) to ensure that (9) does not result in an empty set (infeasibility). This can is achieved with a sufficiently large value of $M$. ∎
III-D Realizable Composite Power Trajectories
The following theorem provides a linear formulation for realizable dispatch of a composite battery.
Theorem 1.
If $\max_{i,j}\{|E^{i}_{0}-E^{j}_{0}|\}\leq\varepsilon$, then any feasible composite battery dispatch schedule $\{P_{\text{c}}[k],P_{\text{d}}[k]\}_{k=0}^{K-1}$ satisfying
[…]
$\frac{P_{\text{c}}[k]}{N{P}_{\text{c,max}}}+\frac{P_{\text{d}}[k]}{N{P}_{\text{d,max}}}\leq\frac{N-1}{N},$ $k=0,\ldots,K-1$ (13c)
$N\varepsilon\leq E[k]\leq N\left(E_{\text{max}}-\varepsilon\right),$ $k=0,\ldots,K$ (13d)
$P_{\text{c}}[k],P_{\text{d}}[k]\geq 0,$ $k=0,\ldots,K-1$ (13e)
will satisfy (1) for all $N$ elements under PSC and is, thus, realizable.
Proof.
Eqs. (1a) and (2) imply (3a) and (1b) implies (3b). From Proposition 1, (13c) ensures satisfaction of (1e). Similarly, with $\max_{i,j}\{|E^{i}_{0}-E^{j}_{0}|\}\leq\varepsilon$, Proposition 2 guarantees that (1c) holds, and the PSC with (13e) and (2) ensure that (1d) holds for all $N$ elements. Thus, any feasible composite dispatch is realizable. ∎
Note that for a fixed composite battery, if $N$ increases, the cutting plane in (13c) approaches the cutting plane in (4b). Furthermore, for $M\gg 1$, $\varepsilon$ in (10) tends to zero. That is, with faster control and more granular elements, (13) approaches the relaxed battery model defined in (4); however, with guaranteed realizability.
The energy buffer $\varepsilon$ in (13d) restricts the set of feasible energy trajectories relative to (3) and (4a). The feasible energy trajectories in the robust linear model presented in [RobustNawaf] are also a subset of the feasible energy trajectories in (3) and (4a). However, the proposed model in (13) has a constant energy buffer $\varepsilon$ that is independent of the length of time horizon considered. This is in contrast to the robust model which has an energy buffer that increases with the length of the time horizon which, in the worst case, significantly reduces the feasible set of energy trajectories and can yield sub-optimal solutions. This is illustrated in Fig. 4 where the (worst case) feasible set of energy trajectories for the robust model and the RCB model for $\Delta t=1\text{hr}$ and $M=1$ ($\delta t=1\text{hr}$) and $M=10$ ($\delta t=6\text{min}$) are compared for the battery parameters of a Tesla Powerwall shown in Table I.Clearly, smaller control time steps result in a smaller $\varepsilon$ and thus a larger feasible set of energy trajectories for the proposed model, while the robust model remains unchanged.
Finally, we note that the PSC policy is a greedy sufficient policy that fully activates BESS elements in sequence. More elaborate policies (greedy or not) can be developed that preferentially spread element loading for most aggregate dispatch signals, without violating the conditions of Theorem 1.
TABLE I: Individual BESS Element Parameters
Parameters Value Unit
Charge/discharge efficiencies, $\eta_{\text{c}}/\eta_{\text{d}}$ 0.95 -
Maximum charge/discharge power, $P_{\text{max}}$ 5.00 kW
Maximum energy capacity, $E_{\text{max}}$ 13.5 kWh
Figure 4: Feasible set energy trajectories for the relaxed model, the robust model and the proposed realizable composite battery model (RCB).
IV Simulation Results
We apply the realizable composite battery (RCB) model in (13) to two relevant use cases and compare its performance against the tighter relaxed model given in (4) and the robust model in [RobustNawaf] used under balanced operation. Also, a Mixed-Integer Linear Program (MILP) formulation with unequal power-sharing for $N$ elements in (1) is used for comparison. The battery element parameters given in Table I were used. The two use-cases are a) power reference tracking and b) revenue maximization.
IV-A Power Reference Tracking
In this use case, the goal is to minimize the power tracking error for a given reference signal as shown below,
$\min_{P_{\text{c}}[k],P_{\text{d}}[k]}\qquad\sum_{k=0}^{K-1}\left((P_{\text{c}}[k]-P_{\text{d}}[k])-P_{\text{ref}}[k]\right)^{2}.$ (14)
A timestep of $\Delta t=5$ min, along with $N=100$ was used. The RCB model was used three times with $M=1$, $M=5$, and $M=10$ respectively. The results from using each model are shown in Fig. 5 and the solve times along with the mean squared tracking error are shown in Table II.
Figure 5: Open-loop implementation of power optimal reference tracking power schedules using MILP, relaxed, robust and proposed RCB models.
TABLE II: Power Reference Tracking Performance Metrics
Model Solver Predicted Actual
Time (ms) MSE (kW2) MSE (kW2)
MILP 50,140 0 0
Relaxed 30 0 1841
Robust 34 2.12 2.12
RCB (M = 1) 30 0.72 0.72
RCB (M = 5) 30 0.013 0.013
RCB (M = 10) 30 0 0
The MILP model yielded the optimal solution and was able to track the reference signal with no tracking error. However, the MILP also came with a significant computational burden compared to the rest of the formulations. The other formulations are linear and, as a result, have significantly faster solve times. The relaxed model predicts zero tracking error; however, the solution was unrealizable and caused the battery to saturate around $t=6$ hrs leading to a large realized tracking error. The robust model yields a realizable yet sub-optimal solution. The RCB model shows improved tracking error with increasing values of $M$ but even with $M=1$ the tracking error surpasses the robust model.
IV-B Revenue Maximization
In this use case, the goal is to maximize the predicted revenue from the composite battery system given a forecasted price signal. The objective function is given by,
$\max_{P_{\text{c}}[k],P_{\text{d}}[k]}\qquad\sum_{k=0}^{K}C[k](P_{\text{d}}[k]-P_{\text{c}}[k]),$ (15)
where $C[k]$ is the price at time $k$. The price signal was obtained from the California Independent System Operator (CAISO) day-ahead market prices for 24 hours, plotted in Fig 6.
Figure 6: CAISO day-ahead price forecasts
A timestep of $\Delta t=15$ min, along with $N=100$ was used. The RCB model was tested for $M=$ 1, 5, and 10 which correspond to $\delta t=$ 15 min, 3 min and 1.5 min respectively. An additional case was added with $M=900$ ($\delta t=1s$) to gain insight on model performance as $M\rightarrow\infty.$The results for the different models are shown in Table III.
TABLE III: Revenue Maximization Performance metrics
Model Solver Predicted Actual
Time (ms) Revenue (\$) Revenue (\$)
MILP 1250 2088.97 2088.97
Relaxed 1 2092.80 1940.35
Robust 1 1866.42 1866.42
RCB (M = 1) 1 1411.53 1411.53
RCB (M = 5) 1 1955.76 1955.76
RCB (M = 10) 1 2022.87 2022.87
RCB (M = 900) 1 2089.23 2089.23
Similar to the previous use-case, the MILP model produced the highest revenue but at a significantly larger computational cost. The relaxed model predicts the highest revenue but underestimates the SOC, leading to unpredicted battery saturation and a realized performance that is significantly worse. The robust model yields a realizable yet suboptimal solution. The RCB model for small values of $M$ results in a sub-optimal solution which can be attributed to the larger timestep used in this use-case leading to a larger value or $\varepsilon$ and, in turn, a more restricted feasible set. However, for f $M\geq 5$ ($\delta t\leq 3$ min), the RCB model surpasses all other linear models and approaches the predicted performance for larger values of $M$. It was also observed that the RCB surpasses the robust model for $M\geq 3$ ($\delta t\leq 5$ min).
V Conclusion
A new linear dispatch model is presented and analyzed for a composite battery that permits unequal power sharing among its $N$ elements to realize simultaneous charging and discharging. A priority-based disaggregation strategy is described and used to characterize sufficient conditions on the composite battery dispatch under which element constraints are guaranteed to be satisfied. Previous battery optimization implicitly assumed equal power sharing between all elements. By enabling unequal power sharing within a linear formulation, we have expanded the class of battery models that yield realizable power trajectories. Practical implementation of such dispatch strategies could also justify the use of linear models for composite storage systems or aggregate storage representations in energy system optimization models.
[…]